# Future-Proofing Class Incremental Learning

### v1

@@ 3.1 Future-Proof Class Incremental Learning
Table 2: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
@@
Table 3: Performances on ImageNet-Subset with the Learning from Half and Learning From Scratch procedures. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
@@
3.2 Generating future data
The synthetic images dataset of future classes is generated following the procedure initially defined by Sariyildiz et al.[43].
The auxiliary dataset $\mathcal{S}$ is generated using a pre-trained diffusion model $G$ conditioned solely on a textual prompt. As proposed in [43], the default prompt used is $c,d_{c}$ with $c$ the name of the class and $d_{c}$ a simple description of the class. Adding the description of the class into the prompt is intended to ensure that the generated images belong to the targeted class and not to any homonyms: for example “crane” in the ImageNet dataset, which may refer to either the bird or the machine used to move heavy objects. The definition $d_{c}$ also tends to restrain the domain of the generated images even though the domain could also be enforced by adding it to the prompt. To automatically create the prompts, we rely on WordNet [30] and use the lemmas of the synset as $c$ and the definition of the synset as $d_{c}$. For each class, $n$ synthetic images are generated with $n$ equals to the number of images per class in the target dataset. Compared to traditional pre-training and concurrent method [2], our proposed method does not require the costly and fastidious process of collecting, curating, and annotating a dataset of real images.
4 Experiments
4.1 Experimental setup
4.1.1 Datasets
The experiments are conducted on two large scale datasets: CIFAR100 [24] and ImageNet-Subset. CIFAR100 is a dataset containing 60,000 32x32 RGB images evenly divided among 100 classes with 500 training images and 100 testing images per class. ImageNet-Subset is a subset of the high-resolution images dataset ImageNet (ILSVRC 2012) [41]. This subset contains 100 classes with about 1,300 training images and 50 testing images per class. Following Rebuffi et al. [39], the classes are ordered using numpy with the random seed 1993.
4.1.2 Model training
Following common practice for exemplar-free class incremental learning [35, 63, 64], ResNet-18 is used for both CIFAR100 and ImageNet-Subset. The feature extractor is trained solely during the first incremental step and is frozen for the rest of the training procedure. The backbone is trained with SDG using the standard Softmax Cross-Entropy loss and AutoAugment [8] for 160 epochs with a batch size of 128 and a weight decay of $5e^{-4}$. The initial learning rate is set to 0.1 and decrease to zero using a cosine annealing schedule. For the classifier, we use a fully-connected linear layer trained following FeTrIL [35]. The classifier is trained using SGD for 50 epochs at each incremental step, with an initial learning rate of 0.1 and cosine annealing schedule. In the ablation study, we also report results using a Nearest Mean Classifier (NMC). Experiments are conducted using Pytorch [33] and PyCIL [59].
@@ 4.1.3 Synthetic generation
Two diffusion models are used for generating the synthetic samples during training: Stable Diffusion [40] and DALL-E2 [37]. If not specified otherwise, the default diffusion model is Stable Diffusion v1.4  [^1] https://huggingface.co/CompVis/stable-diffusion-v1-4 used with a guidance scale of $2.0$ for 50 steps. Following [43], $c,\;d_{c}$ is used for the prompt. The same datasets of synthetic images are used for all the experiments. Those datasets have been generated prior to incremental learning experiments, at size 512x512 using a fixed seed and contains 500 images per class for CIFAR100 and 1300 images per class for ImageNet-Subset. For CIFAR100, the synthetic images are resized to 32x32 before being fed to the model. For DALL-E2, because the model is used through the official API of OpenAI, it is not possible to control the different hyper-parameters of the model; the default, unknown, values are used for the guidance scale, the random seed and the number of denoising step. A slightly different prompt is used for DALL-E2, “a photo of a $c,\;d_{c}$, as the default one would generate unsatisfactory images for some classes. In order to comply with the moderation policy of the API, few definitions $d_{c}$ have been changed.
@@ 4.1.4 Baselines
As a comparison baseline, a dataset of real images containing general and/or similar classes to the ones in the target dataset is used as auxiliary dataset during the first incremental step. This baseline is adapted from the method proposed by Bellitto et al.[2] and applied to exemplar-free class incremental learning with a frozen feature extractor. For CIFAR100, we use CIFAR10 for the auxiliary dataset as both datasets were created using the same annotation and curation methodology from the TinyImages Dataset [48]. However, while CIFAR10 contains more samples than the datasets generated using our proposed method FPCIL, it is important to note that CIFAR10 contains only 10 classes. For ImageNet-Subset, the auxiliary dataset “ImageNet-Compl.” was created by randomly selecting classes from the complement of ImageNet-Subset in ImageNet. “ImageNet-Compl.” contains the same number of classes as the datasets generated by our proposed method FPCIL and approximatively the same number of samples. This is a strong baseline as both subsets are sampled from the same dataset and, while they do not contain the exact same classes, classes in both datasets are from the same categories and some are highly similar: for example both datasets contain different breeds of dogs or species of insects. We additionally use “ImageNet-Compl. X2” which contains two times more classes than ImageNet-Compl. We also report the performance of “FPCIL-Partial (k%)” which generates a dataset of the same size as FPCIL-Oracle but with only $k$of the classes correctly predicted. For CIFAR100, this auxiliary dataset is built by randomly sampling $k$of the classes from FPCIL-Oracle and the remaining ones from the wrong prediction of GPT-3.5. For ImageNet-Subset, “FPCIL-Partial (0%)” contains synthetic images belonging to the same classes as the ones in ImageNet-Compl. The definition of the all different auxiliary datasets used are summarized in Table 1.
@@ 4.2 Results
Table 4: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures for FeTrIL and Nearest Mean Classifier (NMC). Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods CIFAR100
B50 Inc5 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last
FeTrIL [35] 66.81 58.22 50.50 33.86 40.80 24.65
FeTrIL w/ FPCIL-Oracle 72.81 65.20 68.09 55.43 62.92 49.67
   Improvement in p.p. +6.00 +6.98 +17.59 +21.57 +22.12 +25.02
NMC 67.35 57.08 40.28 23.49 29.55 14.34
NMC w/ FPCIL-Oracle 73.57 65.31 62.13 47.81 57.13 42.53
   Improvement in p.p. +6.22 +8.23 +21.85 +24.32 +27.58 +28.19
Table 5: Performances on CIFAR100 with the Learning From Scratch procedure depending on the ratio of correctly predicted future classes in the complementary dataset used during the initial step. Synthetics samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
@@
Table 6: Performances on ImageNet-Subset with the Learning From Scratch procedure depending on the complementary dataset used during initial step. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
@@
Our proposed approach FPCIL considerably improves the average incremental accuracy of the state-of-the-art method FeTrIL [35], from 5.39 percentage points (p.p.) up to 22.12p.p. on CIFAR10 and from 4.24p.p. up to 33.00p.p. on ImageNet-Subset. This improvement is notably higher for the more difficult Learning From Scratch procedure which is particularly challenging for FeTrIL due to the limited quantity of data available during the first incremental step. Compared to using an auxiliary dataset [2] of real images from classes different from the ones in the target dataset, our method also achieves significantly higher accuracy. On ImageNet-Subset, our proposed method FPCIL even outperforms a two times larger auxiliary dataset containing real images. Moreover, even in the case of only wrong guesses for the future classes, the auxiliary dataset of synthetic images performs on par with concurrent approaches relying on real datasets: for the Learning From Half procedure, the difference of average incremental accuracy between using ImageNet-Compl or its synthetic counterpart is at most 0.29p.p.. Furthermore synthetic datasets can be rapidly generated whereas auxiliary datasets of real data require a long and costly collection, curation, and annotation process. All those points highlight the core of our proposed method: while simply using additional real data for pre-training the feature extractor is beneficial, using data from future classes instead of random ones is better, even if the data used are synthetic.
On CIFAR100, Our more practical approach FPCIL-GPT also reaches a higher accuracy than the baseline method FeTrIL yet slightly lower than the FPCIL-Oracle. On CIFAR100 in the Learning From Scratch procedure with 10 steps of 10 classes (B0 Inc10), the less restricted FPCIL-GPT predicted a total of 150 classes with 53 (35,3%) that actually are among the 90 future classes while the most restricted FPCIL-GPT R2 predicted only 53 classes with 30 (56.6%) among the 90 future classes. All the details regarding future class predictions using GPT are reported in Table 1. The more restricted version FPCIL-GPT R1 and R2, tend to have a higher ratio of correctly predicted future classes but also a lower total number of predicted classes, resulting in a lower accuracy than the less restricted FPCIL-GPT. This emphasizes that both the correctness of the prediction of future classes and the actual quantity of data used for pre-training are important. In the ablation study, we further analyze the impact of the correctness of the prediction of future classes.
@@ 4.3 Ablation study
Table 7: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures depending on the Diffusion Model used for generating the synthetic data: Stable Diffusion with a guidance scale of $2.0$, Stable Diffusion with a guidance scale of $7.5$ or DALL-E2. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
@@
Table 8: Performances on CIFAR100  8a) B50 Inc5 and  8b) B0 Inc10 depending on the backbone used. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
@@
4.3.1 Accurately predicting the future
In Table 5, we compare the performance of our proposed methods combined with FeTrIL on CIFAR100 B0 Inc10 depending on the ratio of correctly predicted future classes in the auxiliary dataset used during the initial step. Experimentally, for a fixed dataset size, the average incremental accuracy decreases with the correctness of the prediction of the future classes. This confirms the advantage of predicting the future classes compared to using random classes which may or may not be related to the ones that will be encountered by the continual learner.
@@ 4.3.2 Architecture of feature extractor
4.3.3 Other classifiers
Additionally to FeTrIL [35], we report in Table 4 the performance of our method combined with a Nearest Mean Classifier (NMC). It is a simple yet effective method [19, 32] for class incremental learning with a fixed feature extractor that does not require memory.

### v2

@@ 4.1.2 Model training
Table 2: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-18 as the backbone. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold and second best is underlined. Results averaged over 3 random runs.
@@
Table 3: Performances on ImageNet-Subset with the Learning from Half and Learning From Scratch procedures. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-18 as the backbone. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold and second best is underlined. Results averaged over 3 random runs.
@@ 4.1.3 Synthetic generation
Two diffusion models are used for generating the synthetic samples during training: Stable Diffusion [41] and DALL-E2 [42]. If not specified otherwise, the default diffusion model is Stable Diffusion v1.4  [^1] https://huggingface.co/CompVis/stable-diffusion-v1-4 used with a guidance scale of $2.0$ for 50 steps. Following [43], $c,\;d_{c}$ is used for the prompt. The same datasets of synthetic images are used for all the experiments. Those datasets have been generated prior to incremental learning experiments, at size 512$\times$512 using a fixed seed and contains 500 images per class for CIFAR100 and 1300 images per class for ImageNet-Subset. For CIFAR100, the synthetic images are resized to 32$\times$32 before being fed to the model. For DALL-E2, because the model is used through the official API of OpenAI, it is not possible to control the different hyper-parameters of the model; the default, unknown, values are used for the guidance scale, the random seed and the number of denoising step. A slightly different prompt is used for DALL-E2, “a photo of a $c,\;d_{c}$, as the default one would generate unsatisfactory images for some classes. In order to comply with the moderation policy of the API, few definitions $d_{c}$ have been changed.
@@ 4.1.4 Baselines
As a comparison baseline, a dataset containing real images from classes similar yet different to the ones in the target dataset is used as auxiliary dataset $\mathcal{S}$ during the first incremental step. The purpose of this baseline is to compare our proposed method with the more standard pre-training approach. It is similar to the method proposed by Bellitto et al. [55] adapted to Exemplar-Free Class-Incremental Learning with a frozen feature extractor. For CIFAR100, we use CIFAR10 for the auxiliary dataset as both datasets were created using the same annotation and curation methodology from the TinyImages Dataset [70]. However, while CIFAR10 contains more samples than the datasets generated using our proposed method FPCIL, it is important to note that CIFAR10 contains only 10 classes. For ImageNet-Subset, the auxiliary dataset “ImageNet-Compl.” was created by randomly selecting classes from the complement of ImageNet-Subset in ImageNet. “ImageNet-Compl.” contains the same number of classes as the datasets generated by our proposed method FPCIL and approximately the same number of samples. This is a strong baseline as both subsets are sampled from the same dataset and, while they do not contain the exact same classes, classes in both datasets are from the same categories and some are highly similar: for example both datasets contain different breeds of dogs or species of insects. We additionally use “ImageNet-Compl. X2” which contains two times more classes than ImageNet-Compl. We also report the performance of “FPCIL-Partial (k%)” which generates a dataset of the same size as FPCIL-Oracle but with only $k$of the classes correctly predicted. For CIFAR100, this auxiliary dataset is built by randomly sampling $k$of the classes from FPCIL-Oracle and the remaining ones from the wrong prediction of GPT-3.5. For ImageNet-Subset, “FPCIL-Partial (0%)” contains synthetic images belonging to the same classes as the ones in ImageNet-Compl. The definition of all the different auxiliary datasets used are summarized in Table 1.
@@ 4.2 Results
Our proposed approach FPCIL considerably improves the average incremental accuracy of the state-of-the-art method FeTrIL [7], from 5.39 percentage points (p.p.) up to 22.12p.p. on CIFAR10 and from 4.24p.p. up to 33.00p.p. on ImageNet-Subset. This improvement is notably higher for the more difficult Learning From Scratch procedure which is particularly challenging for FeTrIL due to the limited quantity of data available during the first incremental step. Compared to using an auxiliary dataset of real images from classes different from those in the target dataset, our method also achieves significantly higher accuracy. This highlights the fundamental advantage of our proposed method: while the standard pre-training procedure improves the generalization capability of the feature extractor by leveraging additional data belonging to classes different from those in the target dataset, our proposed method aims to learn a feature extractor specifically designed for the incoming task by relying on data from future classes instead. Even if those data from future classes are synthetic, it allows to learn a feature extractor more suitable for the future tasks resulting in a significantly higher accuracy compared standard pre-training on real data. Moreover, on ImageNet-Subset, our proposed method FPCIL even outperforms a two times larger auxiliary dataset of real images, demonstrating that our proposed algorithm is more data-efficient that standard pre-training. Furthermore synthetic datasets can be rapidly generated whereas auxiliary datasets of real data require a long and costly collection, curation, and annotation process. Finally, even in the case where the future classes have not been accurately predicted, the auxiliary dataset of synthetic images performs on par with the standard pre-training approach relying on real datasets: for the Learning From Half procedure, the difference of average incremental accuracy between using ImageNet-Compl or its synthetic counterpart is at most 0.29p.p.. This makes our proposed method beneficial in every setting.
On CIFAR100, our more practical approach FPCIL-GPT also reaches a higher accuracy than the baseline method FeTrIL yet slightly lower than the FPCIL-Oracle. On CIFAR100 in the Learning From Scratch procedure with 10 steps of 10 classes (B0 Inc10), the less restricted FPCIL-GPT predicted a total of 150 classes with 53 that actually are among the 90 future classes while the most restricted FPCIL-GPT R2 predicted only 53 classes with 30 among the 90 future classes. All the details regarding future class predictions using GPT are reported in Table 1. While the more restricted version FPCIL-GPT R1 and R2 tend to predict a lower number of classes that are not among the future classes, the less restricted version predicts a higher number of future classes, achieving higher performances. There is a trade-off between the precision and recall in the prediction of future classes, while predicting more classes improve the performance it also increase the computational cost of the first incremental step due to the cost of generating synthetic images and training the feature extractor. In the ablation study, we further analyze the impact of the correctness of the prediction of future classes.
Table 4: Performances on CIFAR100 with the Learning From Scratch procedure depending on the ratio of correctly predicted future classes in the complementary dataset used during the initial step. Synthetics samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-18 as the backbone. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold. Results averaged over 3 random runs.
@@
Table 5: Performances on ImageNet-Subset with the Learning From Scratch procedure depending on the complementary dataset used during initial step. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-18 as the backbone. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold. Results averaged over 3 random runs.
@@ 4.3 Ablation study
4.3.1 Accurately predicting the future
In Table 4, we compare the performance of our proposed methods combined with FeTrIL on CIFAR100 B0 Inc10 depending on the ratio of correctly predicted future classes in the auxiliary dataset used during the initial step. Experimentally, for a fixed dataset size, the average incremental accuracy increase with the correctness of the prediction of the future classes. This confirms the importance of predicting the future classes compared to the standard pre-training procedure which relies on real data of classes different from those in the target dataset that will be encountered by the continual learner in the future. Moreover, even in the case where no future class has been correctly predicted, our proposed method still improves the performance of the baseline approach as, in most cases, additional data help to learn a feature extractor with better generalization capability.
4.3.2 Quality and diversity of synthetic samples
Table 6: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures depending on the Diffusion Model used for generating the synthetic data: Stable Diffusion with a guidance scale of $2.0$, Stable Diffusion with a guidance scale of $7.5$, and DALL-E2. Using ResNet-18 as the backbone. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold. Results averaged over 3 random runs.
@@ 4.3.2 Quality and diversity of synthetic samples
Table 7: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures for FeTrIL depending on the number of synthetic samples $n$ generated for each future class. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-18 as the backbone. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods CIFAR100
B50 Inc5 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last
FeTrIL [7] 66.81 58.22 50.50 33.86 40.80 24.65
FeTrIL w/ FPCIL-Oracle $n=50$ 68.31 59.67 60.45 45.56 55.52 40.89
FeTrIL w/ FPCIL-Oracle $n=250$ 71.27 63.23 65.53 52.40 61.45 48.38
FeTrIL w/ FPCIL-Oracle $n=500$ 72.81 65.20 68.09 55.43 62.92 49.67
FeTrIL w/ FPCIL-Oracle $n=1000$ 73.95 66.77 70.24 58.37 64.89 52.73
In Table 7, the performance of FPCIL Oracle is reported on CIFAR100 for different number of synthetic samples $n$ generated for each future classes. While the default value of $n$ equals 500 was used in the other experiments in order to match the size of the target dataset, it appears that just 50 samples per class is enough to improve the performance of the baseline and even outperform the use of eternal dataset of unrelated classes such as CIFAR10 which is more than 10 times bigger. Moreover, it is possible to further improve the performance of our proposed method by generating more synthetic data if the constraints of the system permits it.
In addition to the trade-of between quality and diversity, it is important to emphasize that a significant gap remains between real images and the synthetic samples generated using pre-trained diffusion models. As illustrated in Table 5, while Aux. ImageNet-Compl. and FPCIL-Partial (0%) contain the same classes and almost the same number of samples, there is a difference of about 2 percentage points of accuracy between the feature extractor trained on those datasets. The exact same observation can be made for Aux. ImageNet-Compl. Animals and FPCIL-Partial Animals (0%). Similarly, while achieving significantly higher results than the baseline, FPCIL-Oracle still achieves lower performances on both CIFAR100 and ImageNet-Subset compared to a upper-bound model which would have been trained on the complete target dataset at once.
Table 8: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures for FeTrIL depending on the nature (real or synthetic) of the data used to train the feature extractor during the first incremental step. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-18 as the backbone. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold. Results averaged over 3 random runs.
Methods CIFAR100
Initial classes Future classes B50 Inc5 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last
FeTrIL [7] real - 66.81 58.22 50.50 33.86 40.80 24.65
FeTrIL w/ FPCIL-Oracle var1 - synth. 47.00 44.10 64.05 53.30 64.07 52.44
FeTrIL w/ FPCIL-Oracle var2 synth. synth. 58.04 53.52 65.56 54.42 65.24 53.51
FeTrIL w/ FPCIL-Oracle real synth. 72.81 65.20 68.09 55.43 62.92 49.67
Table 9: Performances on CIFAR100  LABEL:sub@tab:backbone:lfh B50 Inc5 and  LABEL:sub@tab:backbone:lfs B0 Inc10 depending on the backbone used. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold. Results averaged over 3 random runs.
@@
4.3.3 Combining real and synthetic data
During the first incremental step, our proposed method FPCIL combines the real samples of the initial classes with synthetic samples of possible future classes to train the feature extractor. In Table 8, different variations of our proposed method are compared on CIFAR100 to estimate the contribution of each component. “FPCIL var1” does not use the real samples from the first classes and trains the feature extractor solely using synthetic samples of future classes. “FPCIL var2” also generates synthetic samples of the initial classes and train the feature extractor using synthetic samples of both initial and future classes. In both cases, the classifier is then retrained at the end of the first incremental step using the real samples of the initial classes. Our proposed approach achieves the best performance overall, confirming the importance of using the available real data and not relying only on synthetic data. Interestingly, when the number of real data in the first incremental step is extremely small, replacing them by synthetic samples improves the overall performance while decreasing the final accuracy on the first classes. This may be due to bias induced in the learned representation by the distribution gap between real and synthetic samples being exacerbated in the most extreme settings.
@@ 4.3.4 Architecture of feature extractor
4.3.5 Pre-trained model and finetuning
Table 10: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures for FeTrIL using a ResNet-18 feature extractor pre-trained on ImageNet. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-18 as the backbone. Input data are upscaled before being fed to the model in order to match the pre-training configuration. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold. Results averaged over 3 random runs.
Methods CIFAR100
B50 Inc5 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last
FeTrIL [7] 66.73 57.98 51.05 34.43 42.71 26.19
FeTrIL w/ Pre-training 64.16 59.85 71.05 60.97 70.70 59.81
FeTrIL w/ Pre-training + Finetune 76.68 70.55 66.50 54.73 63.05 51.77
FeTrIL w/ FPCIL-Oracle 73.59 65.83 68.10 55.20 63.38 50.54
FeTrIL w/ Pre-training + FPCIL-Oracle 79.06 73.87 76.60 67.14 74.83 64.30
In addition to the main experiments in which the feature extractor is trained from scratch, we also conducted experiments with a pre-trained feature extractor. Using a feature extractor pre-trained on a large external dataset is an effective approach when the low number of classes in the initial incremental step is not enough to train the feature extractor from scratch.
Compared to experiments where the feature extractor is trained from scratch, when using a pre-trained model the initial learning rate for the feature extractor is set to 0.001 and the number of epochs is reduced to 80, other hyper-parameters are kept unchanged. In Table 10, the performances of FeTrIL on CIFAR100 using both a ResNet-18 feature extractor trained from scratch during the first incremental step and a ResNet-18 pre-trained on ImageNet. While it appears that slightly finetuning the feature extractor using the data from the initial incremental step is detrimental if the number of classes in too small, it may also further improve the performance when enough classes are available in the first incremental step, especially if there is a large domain gap between the target dataset and the pre-training dataset. In comparison, FPCIL can be combined with pre-training independently of the number of classes in the first incremental step, increasing the average incremental accuracy by 2.38p.p. up to 5.55p.p..This highlights the versatility of our proposed method FPCIL which can be used either for training the feature extractor from scratch or for finetuning a feature extractor pre-trained on a large dataset, significantly improving the performance of the baseline in both cases.
4.3.6 Comparison with additional baselines
Table 11: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures for FeTrIL, FeCAM, and Nearest Class Mean (NCM) Classifier. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-18 as the backbone. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold. Results averaged over 3 random runs.
Methods CIFAR100
B50 Inc5 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last
NCM 65.24 55.23 40.48 23.37 30.41 15.00
NCM w/ FPCIL-Oracle 71.61 63.21 60.98 46.39 55.69 41.10
Improvement in p.p. +6.37 +7.98 +20.50 +23.02 +25.28 +26.10
FeTrIL [7] 66.81 58.22 50.50 33.86 40.80 24.65
FeTrIL w/ FPCIL-Oracle 72.81 65.20 68.09 55.43 62.92 49.67
Improvement in p.p. +6.00 +6.98 +17.59 +21.57 +22.12 +25.02
FeCAM [33] 70.79 63.04 52.85 38.47 43.30 28.88
FeCAM w/ FPCIL-Oracle 76.77 70.49 71.14 59.95 68.31 56.05
Improvement in p.p. +5.98 +7.45 +18.29 +21.48 +25.01 +27.17
Table 12: Performances on CIFAR100 with the Learning From Scratch procedures. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ Using ResNet-32 as the backbone for CIFAR100 and ResNet-18 for ImageNet-Subset. Results marked with $\dagger$ are reported from [71]. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Best result is marked in bold. Results averaged over 3 random runs.
Methods CIFAR100 ImageNet-Subset
B0 Inc10 B0 Inc5 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last AVG Last
[…]
FeTrIL [7] 46.32 29.80 37.75 21.26 51.64 34.53 39.64 22.42
FeTrIL w/ FPCIL-Oracle 62.73 48.96 59.89 46.62 74.04 65.65 72.64 64.29
FeCAM [33] 44.12 27.83 35.57 20.04 55.37 40.03 43.48 29.11
FeCAM w/ FPCIL-Oracle 64.47 52.01 63.25 50.00 77.78 69.85 77.41 69.17